# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

### v1

@@ 3 Methods
Across our experiments, the original labels are re-scaled to a resolution of $100\;\text{m}\times 100\;\text{m}$, by labeling a patch with the binary class fire if any pixel, within the patch, is labeled as burned. As a consequence, for CanadaFireSat, wildfire forecasting is framed as a dense binary classification task (fire vs. no fire) at the patch level that we name binary patch classification. Moreover, as satellite image time series are not evenly spaced due to cloud cover, we add as complementary information the day of the year for all our predictors. Other details on the experimental setup for all architectures can be found in B. Moreover, we analyze the impact of satellite image time series on model performance in C.
@@ 3.1 CNN-based Architecture
In the CNN-based architecture, satellite image time series are processed in a factorized manner: first spatially and then temporally, for both settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, as shown in Figure 12 and Figure 6, respectively. For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T},\;x_{t}\in\mathbb{R}^{H\times W\times C}$ with $C$ the number bands and the day of the year, and $T$ a fixed number of time steps, each image $x_{t}$ is first encoded independently by a ResNet-50 pre-trained on ImageNet (He et al., 2016): $f(x_{t})=\{z_{i,t}\}_{i=1}^{N_{S}},\;z_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}}$, which outputs $N_{S}=3$ feature maps of channel dimension $D_{i}$, each feature map corresponds to a different scale. The encoding of all time steps is done in parallel, and each scale-specific feature map: $z_{i,t}$, is concatenated independently for each scale across the temporal axis: $z_{i,1:T}=\{z_{i,t}\}_{t=1}^{T}$. Then, the spatio-temporal encoding is done via one ConvLSTM model per scale. By extracting the last hidden state from each ConvLSTM: $g_{i}$, we obtain feature maps $g_{i}(z_{i,1:T})=s_{i},\;s_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times D^{\prime}_{i}}$ at 3 different scales with channel dimension $D^{\prime}_{i}<D_{i}$, providing multiple levels of contextual information.
@@
In setting [Uncaptioned image]satellite images only, our final multi-scale feature maps $\{s_{i}\}_{i=1}^{N_{S}}$ are passed to a U-Net-like decoder. The output of the decoder is interpolated to the dimensions of the label feature map: $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ to match the patch resolution. This is finally passed to binary patch classification layer to output the class probabilities: $h({\{s_{i}\}_{i=1}^{N_{S}}})=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h$ the function representing the decoder, interpolation, and patch classification layer.
@@
In the multi-modal setting [Uncaptioned image]satellite and environmental data, displayed in Figure 6, the above architecture is extended to process in parallel the time series of non-spatial environmental predictors $x_{\text{env},1:T_{\text{env}}}=\{x_{\text{env},t}\}_{t=1}^{T_{\text{env}}},\;x_{\text{env},t}\in\mathbb{R}^{N_{\text{env}}}$ with $N_{\text{env}}$ the number of environmental predictors and $T_{\text{env}}$ a fixed number of time steps. Following (Gorishniy et al., 2022) for tabular data encoding, each environmental predictor is projected to a high-dimension space through specific MLP layers: $\forall j\in\{1,\dots,N_{\text{env}}\},\;f_{\text{env},j}(x_{\text{env},t}^{j})=z_{\text{env},t}^{j},\;z_{\text{env},t}^{j}\in\mathbb{R}^{D_{\text{env}}}$. The projected features are then averaged across the $N_{\text{env}}$ dimension to obtain $z_{\text{env},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{env}}\times T_{\text{env}}}$ and passed to an LSTM model for temporal encoding $g_{\text{env}}(z_{\text{env},1:T_{\text{env}}})=s_{\text{env}},\;s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ for the final environmental encoded feature. This one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder before the patch classification layer: $h({\{s_{i}\}_{i=1}^{N_{S}}},s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
@@
For setting [Uncaptioned image]environmental predictors only, where the model is trained using only environmental predictors at a resolution varying from $500$ m up to $28$ km (see Table 2), we first split the predictors into two groups: mid-resolution inputs $x_{\text{mid},1:T_{\text{env}}}=\{x_{\text{mid},t}\}_{t=1}^{T_{\text{env}}},\;x_{\text{mid},t}\in\mathbb{R}^{H_{\text{mid}}\times W_{\text{mid}}\times N_{\text{mid}}}$ for all MODIS data, and low-resolution inputs for all ERA5 and CEMS data: $x_{\text{low},1:T_{\text{env}}}=\{x_{\text{low},t}\}_{t=1}^{T_{\text{env}}},\;x_{\text{low},t}\in\mathbb{R}^{H_{\text{low}}\times W_{\text{low}}\times N_{\text{low}}}$. In this setting, we leverage spatial environmental inputs and not tabular to compensate for the absence of high-resolution satellite imagery (Sentinel-2), providing spatial context for settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data. In each group, as not all predictors have the same spatial resolution (see Table 2), we upsampled all predictors to the highest available resolution. Details on the different spatial dimensions for each group can be found in the B. We will partially modify the architecture from setting [Uncaptioned image]satellite and environmental data, as shown in Figure 13. First, mid-resolution inputs are used as an alternative to satellite image time series. In practice, all the satellite image processing model components stay the same for the mid-resolution inputs group: the spatial encoding $f$, the scale-specific temporal encoding $g_{i}$, and the final head $h$ (corresponding to the decoder, interpolation layer, and patch classification layer). We simply extend the number of multi-scale feature maps to $N_{S}=5$ because of the lower resolution of the input data. Moreover, we exchange ConvLSTM with LSTM when the output feature maps from $f$ become one-dimensional (for $i=5$). The second branch of the model is adapted to process spatial data for the low-resolution inputs group. We use a smaller pre-trained CNN architecture to encode independently each time step similarly to satellite images: ResNet-18 (He et al., 2016): $f_{\text{low}}(x_{\text{low},t})=z_{\text{low},t},\;z_{\text{low},t}\in\mathbb{R}^{D_{\text{low}}}$. The spatial encoding for this input resolution outputs a single-scale 1-D vector. The temporally concatenated features $z_{\text{low},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{low}}\times T_{\text{env}}}$ are passed to an LSTM model $g_{\text{low}}(z_{\text{low},1:T_{\text{env}}})=s_{\text{low}},\;s_{\text{low}}\in\mathbb{R}^{D^{\prime}_{\text{low}}}$ to obtain low-resolution encoded features with $D^{\prime}_{\text{low}}<D_{\text{low}}$. Similarly to the multi-modal architecture, this one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder that has processed the mid-resolution group.
@@
In all three settings, the training is done via a per-patch loss: $L_{\text{CNN}}$, which is a combination of weighted cross-entropy loss and dice loss. Weighted cross-entropy gives more importance to the rare class: fire, by increasing its loss contribution, while dice loss measures overlap: IoU, and directly optimizes for better segmentation of small or imbalanced regions. The losses are as follows:
@@
where $y_{i}$ is the ground truth label for a patch (1 for fire, 0 for no fire), $\hat{y}_{i}$ is the predicted probability for fire, and $w_{\text{fire}}$ and $w_{\text{no fire}}$ are the class weights.
@@ 3.2 Transformer-based Architecture
In the three settings, our ViT architectures re-use most of the components of their CNN counterparts, as shown in Figure 7, 14, and 15. The main difference is the absence of multi-scale feature maps after the satellite image encoding in options [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, or after the mid-resolution encoding for option [Uncaptioned image]environmental predictors only.
For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T}$, each image $x_{t}\in\mathbb{R}^{H\times W\times C}$ is encoded independently by a pre-trained ViT architecture, specifically DINOv2: ViT-S (Oquab et al., 2023): $f(x_{t})=\{z_{t}\},\;z_{t}\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$, which outputs one feature map per time-step. Similarly to the CNN architecture, the encoding of all satellite images time steps is done in parallel, and the feature maps are concatenated across the temporal axis: $z_{1:T}=\{z_{t}\}_{t=1}^{T}$. As for the CNN, the temporal encoding is also done via a ConvLSTM model: $g(z_{1:T})=s,\;s\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$. Multi-scale feature maps are not necessary for ViT due to the native high-resolution of the final output feature map: $H_{p}$ and $W_{p}$. The output feature map: $s$, is interpolated to the label dimensions: $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ and finally passed to the model head: a patch classification layer, to output the class probabilities: $h_{\text{ViT}}(s)=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h_{\text{ViT}}$ the function representing the interpolation, and classification layer.
In the multi-modal extension [Uncaptioned image]satellite and environmental data, the encoding of environmental inputs is identical to that of the CNN method. The final environmental encoded features $s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ is replicated spatially and concatenated with the final feature map $s$ before the patch classification layer to output the class probabilities: $h(s,s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
@@
Contrary to the training for CNN-based architecture, the loss used here is only the dice loss as defined in Equation 6, because experimentally it led to the best results.
@@ 4 Results
This section details the key results for the benchmark models described in Section 3. CanadaFireSat covers the period 2016-2023: we train our models on years 2016-2021, while keeping 2022 for validation (Val) and 2023 for both test sets: Test and Test Hard (see Section 4.1 for details). Table 4 reports the performance of our benchmark models on CanadaFireSat. Results are evaluated in terms of F1 score and PRAUC (Area Under the Precision-Recall Curve for the positive class fire only), both being robust to imbalanced datasets. The F1 score is defined as the harmonic mean between Precision (proportion of true positive pixels over pixels predicted as positive) and Recall (proportion of true positives over all actual positives):
@@
It provides information about how well the model minimizes both false negatives and false positives at a fixed threshold. PRAUC summarizes the Precision-Recall trade-off across all thresholds for the class fire.
The benchmark models are tested against a baseline approach relying on the FWI. Wildfire forecasting maps are obtained from the FWI in the following way. First, for a given time step $t$ we extract the 8-day mean FWI map at $0.25$° over Canada from $t-8$ to $t-1$ included, and interpolate it at the target resolution of $100\>m\times 100\>m$. Then, the per-patch prediction is obtained by binarizing the interpolated FWI map. The optimal threshold is tuned on the validation set, referring to the year 2022: $\text{FWI}_{\text{th}}=6$. The PRAUC is computed by scaling the FWI values with respect to the maximum value: $\text{FWI}_{\text{max}}=50$.
4.1 Evaluation on different sets: Val, Test, and Test Hard
To assess the performance of the models, we rely on three different evaluation sets: Val, Test, and Test Hard. The validation (Val) and test sets (Test) represent the years 2022 and 2023 of our CanadaFireSat dataset, respectively. As such, negative examples present in these two sets follow the same sampling scheme as described in Section 2. Following a standard pipeline in machine learning, we tune the model’s hyperparameters and threshold using the F1 score on the Val set, before deploying the models on the Test set. From Table 4, a significant performance gap between Val and Test is observed for all the benchmarked models, with an increase in both PRAUC and F1 score on Test. This shift can be explained by a strong inter-annual variability between the year 2022 (Val) and the year 2023 (Test). Indeed, 2023 was a record-breaking wildfire season in Canada (Jain et al., 2024) ($14.01$ mha in 2023 against $1.38$ in 2022, resulting in $19.6\%$ of fire patches in the Test set compared to $11.5\%$ in the Val set), mainly driven by warm and dry weather. Due to such extreme environmental conditions, 2023 could be considered as a simple test set when assessing the behavior of our model, with the baseline relying on the FWI only already reaching a very high F1 score of $50.3$. To evaluate our models’ performance on a more challenging test set, we created a second test set named Test Hard based on an adversarial negative sampling scheme, where the FWI would be less discriminative to distinguish between the class fire and the class no fire. To make the distribution of negative samples similar to that of positives in terms of fire risk, we perform a stratified sampling of the year 2023 by extending Equation 1 to account for FWI buckets, land cover, and month of the year. Land cover is downloaded from ESA WorldCover at $10\>m$ for 2020. Further analysis of the model performance threshold analysis across the three sets is shown in D.
The resulting distribution is shown in Figure 8. With this new sampling scheme, the negative sample distribution is much more similar to that of positive samples across the FWI. As a result, we can hypothesize that, in Test Hard, ignition acts as the main discriminative factor between the fire and no fire classes. As such, Test Hard represents adversarial examples for a model that learned for wildfire fuel and weather patterns. By deploying the trained networks on Test Hard, we can assess their performance for modeling ignition, a complex triggering factor that is not taken into account by the FWI, but whose pattern can be implicitly learned by deep models from the training data. For this reason, the performance of our trained models for high-resolution wildfire forecasting on Test Hard can be considered as a lower bound for such an extreme wildfire season. Further details about the distribution of samples across land-cover classes are provided in Figure 17.

### v2

@@ 3 Methods
For CanadaFireSat, wildfire forecasting is framed as a binary classification task (fire vs no fire) at the patch level, i.e., a binary patch classification. Across our experiments, the original labels at a native resolution of $10\;\text{m}\times 10\;\text{m}$ are re-scaled to $100\;\text{m}\times 100\;\text{m}$, by labeling a patch with the binary class fire if any pixel within the patch is labeled as burned. This design decision aims to focus on providing alerts for any size of fires at the expense of false positive pixels at the native resolution and is often used in wildfire prediction at both coarse (Prapas et al., 2022; Bakke et al., 2023) and high-resolution (Pelletier et al., 2023). It is also motivated by the shortcomings of MODIS, in particular the MCD64A1 burned area product, which is recurrently used in coarse wildfire forecasting (Huot et al., 2020; Rodrigues et al., 2022; Prapas et al., 2021) despite underestimating burned area (Bakke et al., 2023; Zhu et al., 2017).
Finally, as satellite image time series are not evenly spaced due to cloud cover, we add as complementary information the day of the year for all our predictors composing the time series. Other details on the experimental setup for all architectures can be found in B. We analyze the impact of satellite image time series on model performance in C.
@@ 3.1 CNN-based Architecture
In the CNN-based architecture, satellite image time series are processed in a factorized manner: first spatially and then temporally, for both settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, as shown in Figure 12 and Figure 7, respectively. For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T}$, with $x_{t}\in\mathbb{R}^{H\times W\times C}$ being a single time step with $C$ the number bands and the day of the year, and $T$ a fixed number of time steps, each image $x_{t}$ is first encoded independently by a ResNet-50 pre-trained on ImageNet (He et al., 2016): $f(x_{t})=\{z_{i,t}\}_{i=1}^{N_{S}}$, with, $z_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}}$, which outputs $N_{S}=3$ feature maps of channel dimension $D_{i}$, each feature map corresponding to a different scale. The encoding of all time steps is done in parallel, and each scale-specific feature map, $z_{i,t}$, is concatenated independently for each scale across the temporal axis: $z_{i,1:T}=\{z_{i,t}\}_{t=1}^{T}$. Then, the spatio-temporal encoding is done via one ConvLSTM model per scale. By extracting the last hidden state from each ConvLSTM: $g_{i}$, we obtain feature maps $g_{i}(z_{i,1:T})=s_{i}$, with $s_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times D^{\prime}_{i}}$ at 3 different scales with channel dimension $D^{\prime}_{i}<D_{i}$, providing multiple levels of contextual information.
@@
In setting [Uncaptioned image]satellite images only (Figure 12), our final multi-scale feature maps $\{s_{i}\}_{i=1}^{N_{S}}$ are passed to a U-Net-like decoder. The output of the decoder is interpolated to the dimensions of the label feature map: $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ to match the patch resolution. This is finally passed to binary patch classification layer to output the class probabilities: $h({\{s_{i}\}_{i=1}^{N_{S}}})=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h$ the function representing the decoder, interpolation, and patch classification layer.
@@
In the multi-modal setting [Uncaptioned image]satellite and environmental data (Figure 7), the above architecture is extended to process in parallel the time series of non-spatial environmental predictors $x_{\text{env},1:T_{\text{env}}}=\{x_{\text{env},t}\}_{t=1}^{T_{\text{env}}}$, with $x_{\text{env},t}\in\mathbb{R}^{N_{\text{env}}}$ being the data for a single time step with $N_{\text{env}}$ environmental predictors. $T_{\text{env}}$ is a fixed number of time steps. Following (Gorishniy et al., 2022) for tabular data encoding, each environmental predictor is projected to a high-dimensional space with specific MLP layers: $\forall j\in\{1,\dots,N_{\text{env}}\},\;f_{\text{env},j}(x_{\text{env},t}^{j})=z_{\text{env},t}^{j}$, with $z_{\text{env},t}^{j}\in\mathbb{R}^{D_{\text{env}}}$. The projected features are then averaged across the $N_{\text{env}}$ dimension to obtain $z_{\text{env},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{env}}\times T_{\text{env}}}$ and passed to an LSTM model for temporal encoding $g_{\text{env}}(z_{\text{env},1:T_{\text{env}}})=s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ , which we use as the final environmental encoded feature. This one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder before the patch classification layer: $h({\{s_{i}\}_{i=1}^{N_{S}}},s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
@@
For setting [Uncaptioned image]environmental predictors only (Figure 13), where the model is trained using only environmental predictors at a resolution varying from $500$ m up to $28$ km (see Table 2), we first split the predictors into two groups: mid-resolution inputs $x_{\text{mid},1:T_{\text{env}}}=\{x_{\text{mid},t}\}_{t=1}^{T_{\text{env}}}$, with single time step $x_{\text{mid},t}\in\mathbb{R}^{H_{\text{mid}}\times W_{\text{mid}}\times N_{\text{mid}}}$ for all MODIS data, and low-resolution inputs for all ERA5 and CEMS data: $x_{\text{low},1:T_{\text{env}}}=\{x_{\text{low},t}\}_{t=1}^{T_{\text{env}}}$, with single time step $x_{\text{low},t}\in\mathbb{R}^{H_{\text{low}}\times W_{\text{low}}\times N_{\text{low}}}$. In this setting, we leverage spatial environmental inputs and not tabular to compensate for the absence of high-resolution satellite imagery from Sentinel-2, which provides spatial context for settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data. In each group, as not all predictors have the same spatial resolution (see Table 2), we upsampled all predictors to the highest available resolution. Details on the different spatial dimensions for each group can be found in the B. We partially modify the architecture from setting [Uncaptioned image]satellite and environmental data, as shown in Figure 13. First, mid-resolution inputs are used as an alternative to satellite image time series. In practice, all the satellite image processing model components stay the same for the mid-resolution inputs group: the spatial encoding $f$, the scale-specific temporal encoding $g_{i}$, and the final head $h$ (corresponding to the decoder, interpolation layer, and patch classification layer). We simply extend the number of multi-scale feature maps to $N_{S}=5$ because of the lower resolution of the input data. Moreover, we exchange ConvLSTM with LSTM when the output feature maps from $f$ become one-dimensional (for $i=5$). The second branch of the model is adapted to process spatial data for the low-resolution inputs group. We use a smaller pre-trained CNN architecture to encode independently each time step, similarly to the processing of satellite images described above: we use ResNet-18 (He et al., 2016) to obtain a one-dimensional feature vector $f_{\text{low}}(x_{\text{low},t})=z_{\text{low},t}$, with $z_{\text{low},t}\in\mathbb{R}^{D_{\text{low}}}$. The temporally concatenated features $z_{\text{low},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{low}}\times T_{\text{env}}}$ are passed to an LSTM model $g_{\text{low}}(z_{\text{low},1:T_{\text{env}}})=s_{\text{low}}$, with $s_{\text{low}}\in\mathbb{R}^{D^{\prime}_{\text{low}}}$ to obtain low-resolution encoded features with $D^{\prime}_{\text{low}}<D_{\text{low}}$. Similarly to the multi-modal architecture, this one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder that has processed the mid-resolution group.
@@
In all three settings, the training is done with a per-patch loss, $L_{\text{CNN}}$, which is a combination of weighted cross-entropy loss and dice loss. Weighted cross-entropy gives more importance to the rare class fire by increasing its contribution to the loss, while the dice loss measures overlap (i.e. intersection over union) and directly optimizes for better segmentation of small or imbalanced regions. The losses are as follows:
@@
where $y_{i}$ is the ground truth label for a patch (1 for fire, 0 for no fire), $\hat{y}_{i}$ is the predicted probability for fire, and $w_{\text{fire}}$ and $w_{\text{no fire}}$ are class weights.
@@ 3.2 Transformer-based Architecture
In the three settings, our ViT architectures re-use most of the components of their CNN counterparts, as shown in Figure 8, 14, and 15, respectively. The main difference is the absence of multi-scale feature maps after the satellite image encoding in options [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, or after the mid-resolution encoding for option [Uncaptioned image]environmental predictors only.
For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T}$, each image $x_{t}\in\mathbb{R}^{H\times W\times C}$ is encoded independently by a pre-trained ViT architecture, specifically DINOv2: ViT-S (Oquab et al., 2023): $f(x_{t})=\{z_{t}\}$, with $z_{t}\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$, which outputs one feature map per time-step. Similarly to the CNN architecture, the encoding of all satellite images time steps is done in parallel, and the feature maps are concatenated across the temporal axis: $z_{1:T}=\{z_{t}\}_{t=1}^{T}$. As for the CNN, the temporal encoding is also done via a ConvLSTM model: $g(z_{1:T})=s$, with $s\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$. Multi-scale feature maps are not necessary for ViT due to the native high-resolution of the final output feature map: $H_{p}$ and $W_{p}$. The output feature map, $s$, is interpolated to the label dimensions $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ and finally passed to the model head, a patch classification layer, to output the class probabilities: $h_{\text{ViT}}(s)=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h_{\text{ViT}}$ the function representing the interpolation, and classification layer.
In the multi-modal model ([Uncaptioned image]satellite and environmental data), the encoding of environmental inputs is identical to that of the CNN method. The final environmental encoded features $s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ is replicated spatially and concatenated with the final feature map $s$ before the patch classification layer to output the class probabilities: $h(s,s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
@@
Contrary to the training for the CNN-based architectures, the loss used here is only the dice loss as defined in Equation 7, because experimentally it led to the best results.
@@ 4 Results
This section details the key results for the benchmark models described in Section 3. CanadaFireSat covers the period 2016-2023: we train our models on the years 2016-2021, while keeping 2022 for validation (Val) and 2023 for both test sets: Test and Test Hard. Results are evaluated in terms of F1 score and PRAUC (Area Under the Precision-Recall Curve for the positive class fire only). Both metrics are robust to imbalanced datasets, contrarily to patch-level accuracy. The F1 score is defined as the harmonic mean between Precision (proportion of true positive pixels over pixels predicted as positive) and Recall (proportion of true positives over all actual positives):
@@
It provides information about how well the model minimizes both false negatives and false positives at a fixed threshold. We favor the F1 score over Intersection over Union (IoU) as the former is more commonly used in the wildfire forecasting literature. PRAUC summarizes the Precision-Recall trade-off across all thresholds for the class fire.
The benchmark models are tested against a baseline approach relying on the FWI in the following way: first, for a given time step $t$ we extract the 8-day mean FWI map at 0.250.25° from t8t-8 to t1t-1 included, and interpolate it at the target resolution of $100\>\text{m}\times 100\>\text{m}$, then, the per-patch prediction is obtained by binarizing the interpolated FWI map. The optimal threshold is tuned on the validation set, referring to the year 2022: $\text{FWI}_{\text{th}}=6$. The PRAUC is computed by scaling the FWI values with respect to the maximum value: $\text{FWI}_{\text{max}}=50$.